OpenStudy (anonymous):

use substitution method (1) 0.3x + 0.2y = 5 (2) 0.5x + 0.4y = 11

OpenStudy (callisto):

0.3x + 0.2y = 5 => 0.2y = 5- 0.3x -(1) 0.5x + 0.4y = 11 => 0.5x + 2(0.2y) =11 -(2) Sub. (1) into (2) 0.5x + 2( 5- 0.3x ) =11 0.5x + 10 - 0.6x =11 -0.1x = 1 x=?

OpenStudy (anonymous):

Callisto thankyou so much. I was a lil close!!!

OpenStudy (callisto):

Welcome!~